OpenStudy (anonymous):

Determine the zeros of f(x) = x3 – 12x2 + 28x – 9. (1 point)

OpenStudy (anonymous):

x²(x-12)+ i wasnt sure how to do the next part because nothing in goes into 9 and 28 . i dont think .

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

since its not working we need to do it lil different

ganeshie8 (ganeshie8):

lets split the given polynomial into 6 terms :-

ganeshie8 (ganeshie8):

\(f(x) = x^3 – 12x^2 + 28x – 9 \) \( = x^3 – 9x^2 - 3x^2 + 27x + x – 9 \)

ganeshie8 (ganeshie8):

i just wrote, -12x2 as -9x2 - 3x2 28x as 27x + x

ganeshie8 (ganeshie8):

is that fine ?

OpenStudy (anonymous):

yes . okk i see what you did there.

ganeshie8 (ganeshie8):

next, first two terms pull out x^2

ganeshie8 (ganeshie8):

middle two terms pull out -3x

ganeshie8 (ganeshie8):

last two terms pull out 1

ganeshie8 (ganeshie8):

\(f(x) = x^3 – 12x^2 + 28x – 9 \) \( = x^3 – 9x^2 - 3x^2 + 27x + x – 9 \) \( = x^2(x – 9) - 3x(x - 9) + 1(x – 9) \)

ganeshie8 (ganeshie8):

now (x-9) is common, pull it out

OpenStudy (anonymous):

x3 – 9x²-3x²+27x+x-9 x²(x-9)-3x(x-9)+1(x-9)

OpenStudy (anonymous):

okk thats what i got.

OpenStudy (anonymous):

okk so (x²-3x+1)(x-9)

ganeshie8 (ganeshie8):

\(f(x) = x^3 – 12x^2 + 28x – 9 \) \( = x^3 – 9x^2 - 3x^2 + 27x + x – 9 \) \( = x^2(x – 9) - 3x(x - 9) + 1(x – 9) \) \( =(x – 9)( x^2 - 3x + 1)\)

ganeshie8 (ganeshie8):

yes, thats the maximum we can factor. now set it to 0

ganeshie8 (ganeshie8):

\((x – 9) (x^2-3x+1) = 0\)

OpenStudy (anonymous):

so then x-9=0 and x²-3x+1=0

ganeshie8 (ganeshie8):

yes, you need to solve the quadratic

OpenStudy (anonymous):

so x=9. im not sure how to do the other one :(

ganeshie8 (ganeshie8):

thats one zero, x = 9

ganeshie8 (ganeshie8):

x²-3x+1=0

ganeshie8 (ganeshie8):

use quadratic formula

ganeshie8 (ganeshie8):

\(\large x = \frac{-b \pm \sqrt{b^2-4ac}}{2a}\)

ganeshie8 (ganeshie8):

\(\large x = \frac{-(-3) \pm \sqrt{(-3)^2-4 \times 1 \times 1}}{2 \times 1}\)

ganeshie8 (ganeshie8):

\(\large x = \frac{3 \pm \sqrt{9-4}}{2}\)

ganeshie8 (ganeshie8):

\(\large x = \frac{3 \pm \sqrt{5}}{2}\)

OpenStudy (anonymous):

okk sooo im there so far. then idont what to do from there.

ganeshie8 (ganeshie8):

yes thats where we stop :)

ganeshie8 (ganeshie8):

so, the zeroes are : \(9, \frac{3 + \sqrt{5}}{2}, \frac{3 - \sqrt{5}}{2}\)

OpenStudy (anonymous):

ohhhh lol wooww thank you soo much . you are amazing !

ganeshie8 (ganeshie8):

ohh ty :) you're amazingly brilliant too :D